OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $7.65. If the number of dimes was doubled and the number of nickels was increased by 35, the value of the coins would be $14.30. How many nickels and dimes does he have? (Points : 20)

OpenStudy (anonymous):

let number of dimes = x and number of nickles = y 1 dime = 10 cents 1 nickle = 5 cents value of dimes = 10*x cents = 10x cents value of nickles = 5*y cents = 5y cents they totaled to $7.65 = 765 cents so 10x+5y=765 ----eqn(I) number of dimes was doubled. so number of dimes = 2x number of nickles was increased by 35 so number of nickles = y+35 value of dimes = 10*2x cents = 20x cents value of nickles = 5*(y+35) cents = (5y + 175) cents they totaled to $14.30=1430 cents so 20x + 5y +175 = 1430 so 20x + 5y = 1255 ----eqn(II) solve eqn(I) and (II) to get values of x=49 and y=55 so number of dimes = x = 49 and number of nickles = y =55